# Personalized Item Representations in Federated Multimodal Recommendation

### v2

@@ 5.6.3. Ablation Study on Fusion Strategy
5.6.4. Privacy Protection
Table 4 presents the experimental results of the used FedRec methods combined with FedMR, comparing the performance with and without adding Gaussian noise (mean=0, variance=0.1) to the gradients of the global components uploaded to the server, to protect user privacy. The experiments are conducted on the KU dataset during feature fusion. The results indicate that all methods experienced some performance degradation after adding noise, particularly on the NDCG metric, which is more sensitive to ranking accuracy. This is because noise disturbs the gradient precision, which affects the model’s ability to update effectively.
Among the methods, FedNCF is the most sensitive to noise, with HR@50 dropping by 14.33% and NDCG@50 by 31.03%. As a complex neural network-based model, FedNCF depends on numerous parameters to capture nonlinear features. The introduction of noise disrupts these feature interactions, resulting in a significant decline in performance. In contrast, FedAvg and FedRAP show greater robustness to noise, with only minimal performance degradation. This suggests that these methods are more effective at maintaining recommendation accuracy while ensuring privacy protection.
In conclusion, although adding noise helps protect user privacy, its impact varies based on model complexity and feature interaction mechanisms. Simpler models like FedAvg and adaptive models like FedRAP are more resilient to noise, while complex neural network models like FedNCF experience more significant performance drops due to the noise interference.
@@ 6. Conclusion
This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client. This module dynamically adjusts fusion weights based on user interaction history, generating personalized item representations that capture fine-grained user preferences, thereby enhancing personalized recommendations. It allows flexible modification of fusion strategies and supports different modalities of item data as needed. Experimental results on four multimodal recommendation datasets show that FedMR performs well and is compatible with existing ID-based methods without requiring changes to the underlying framework. This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.
@@ A.3. Global Aggregation
Appendix B Parameter Analysis
@@
Despite the increase in the number of learnable parameters, Table 2 demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset. By combining multimodal features with ID embeddings, FedMR better leverages users’ interaction histories to capture the fine-grained personality of users. Additionally, the mixing feature fusion module in FedMR adjusts the weights of multimodal features based on users’ interaction history, greatly enhancing both the accuracy and personalization of recommendations. Moreover, compared to traditional FedRec methods, which rely solely on the ID embeddings, FedMR enables personalized modeling by integrating multimodal information, thereby further improving recommendation quality. While the model’s complexity increases, this trade-off results in significant performance gains.
@@
Appendix C More Experiments
@@
C.1. Ablation Study on Client Sampling
@@
Appendix D Limitations and Future Work
While FedMR has shown promising improvements in federated multimodal recommendation systems, several limitations remain:
  1. (1)
    Data Heterogeneity: Considering the federated settings, the distribution of user data varies significantly. Although FedMR mitigates this issue by integrating multimodal information, its performance is still limited when handling highly heterogeneous data. Future work could focus on introducing more robust personalization models or adaptive aggregation strategies to better address the differences between users.
  2. (2)
    Communication Overhead: Although FedMR reduces some communication costs by encoding multimodal data, processing multimodal features still results in high computational and communication costs. Future research could explore optimizing communication protocols or using model compression techniques to further minimize these costs.
  3. (3)
    Limited Flexibility in Multimodal Fusion: While FedMR supports dynamic adjustments to its fusion strategies, its current approach may lack flexibility when dealing with complex user behaviors. Future work could explore more adaptive and layered fusion methods to better capture the diverse preferences of users.
In conclusion, future developments could enhance the overall performance of federated recommendation systems by improving the handling of data heterogeneity, reducing communication costs, and optimizing multimodal fusion strategies.

### v3

@@ 5.6.3. Ablation Study on Client Heterogeneity
5.6.4. Ablation Study on Privacy Preservation
A key consideration in FedRec is compatibility with privacy-enhancing technologies. We assess FedMR’s robustness by evaluating the performance impact of adding Gaussian noise to client gradients before server aggregation, a common technique related to differential privacy. Table 4 presents these results on the KU dataset. As anticipated, introducing noise creates a privacy-utility trade-off, generally leading to decreased recommendation accuracy across all tested FedMR-enhanced models. However, the extent of this performance degradation differs significantly depending on the base FedRec architecture. Models with simpler structures or strong adaptive components, specifically FedAvg, FCF, and FedRAP, demonstrate considerable resilience, experiencing relatively minor performance reductions. Conversely, methods employing more complex neural networks, such as FedNCF, exhibit higher sensitivity to the injected noise, resulting in a more substantial drop in accuracy. This differential impact suggests that noise may interfere more significantly with the intricate parameter relationships learned by complex models. Nevertheless, the performance reduction remains moderate for most configurations, indicating that FedMR effectively integrates with various baseline methods and maintains reasonable utility under standard noise-based privacy mechanisms, confirming its practical applicability for privacy-conscious recommendation tasks in federated environments.
@@ 6. Conclusion
Motivated by privacy concerns and regulations that necessitate alternatives to data-intensive centralized multimodal recommendation, this paper introduced FedMR, a novel framework addressing the critical challenge of effectively integrating multimodal data within privacy-preserving Federated Recommendation Systems. FedMR’s core innovation lies in the client-side Mixing Feature Fusion Module, which facilitates dynamic, personalized adaptation by fusing server-provided multimodal representations with ID-based signals according to individual user interaction histories, explicitly modeling and addressing user heterogeneity in multimodal preference fusion directly on the client. This adaptive mechanism effectively handles diverse user preferences across modalities. Extensive experiments demonstrate that FedMR seamlessly enhances various ID-based FedRecs, substantially boosting their performance and significantly narrowing the gap towards centralized multimodal systems, occasionally even surpassing them. By enabling adaptive multimodal fusion under federated constraints, FedMR offers a practical and robust solution, advancing the state-of-the-art towards effective, personalized, and privacy-preserving multimodal recommendation in federated environments.
@@ A.3. Global Aggregation
Appendix B Theoretical Analysis
B.1. Convergence Analysis
Proof of Lemma 1 (Local Update Bound).
Using the local update rule:
(16) $\Theta_{l,u}^{t,k}=\Theta_{l,u}^{t,k-1}-\eta\,g_{l,u}(\Theta_{l,u}^{t,k-1};\xi),$
and Assumption 2, we have:
(17) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,k}-\Theta_{l,u}^{t}\Bigr{\|}^{2}&=\mathbb{E}\left\|\sum_{i=1}^{k}\eta\,g_{l,u}(\Theta_{l,u}^{t,i-1};\xi)\right\|^{2}\\[4.0pt] &\leq k\,\eta^{2}\sum_{i=1}^{k}\mathbb{E}\left\|g_{l,u}(\Theta_{l,u}^{t,i-1};\xi)\right\|^{2}\quad\text{(by Jensen's inequality)}\\[4.0pt] &\leq k^{2}\,\eta^{2}(G^{2}+\sigma_{l}^{2})\quad\text{(by Assumption 2).}\end{split}$
Jensen’s inequality is used because the square of a sum under an expectation can be bounded by $k$ times the sum of squares.
For $\eta=\tfrac{1}{L\sqrt{A}}$ and $k=A$, it follows that:
(18) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}&\leq A^{2}\,\eta^{2}(G^{2}+\sigma_{l}^{2})=\frac{A(G^{2}+\sigma_{l}^{2})}{L^{2}}.\end{split}$
Defining $C_{1}=A$, we can equivalently express the bound as the followings:
(19) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}\leq C_{1}\,A\,\eta^{2}(G^{2}+\sigma_{l}^{2}).\end{split}$
Proof of Lemma 2 (Gradient Bias Bound).
The gradient bias is given by
(20) $\begin{split}\|B_{t}\|&=\Bigl{\|}\mathbb{E}_{u}\Bigl{[}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}\Bigr{]}-\nabla_{\Theta_{g}}F(\Theta_{g}^{t})\Bigr{\|}\\[4.0pt] &=\Bigl{\|}\mathbb{E}_{u}\Bigl{[}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}-\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t}\bigr{)}\Bigr{]}\Bigr{\|}\\[4.0pt] &\leq\mathbb{E}_{u}\Bigl{\|}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}-\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t}\bigr{)}\Bigr{\|}\\[4.0pt] &\leq L\,\mathbb{E}_{u}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|},\end{split}$
where the last inequality follows from Assumption 1. Squaring both sides and applying the result from Lemma 1 yields:
(21) $\begin{split}\|B_{t}\|^{2}\leq L^{2}\,\mathbb{E}_{u}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}\leq L^{2}\,C_{1}\,A\,\eta^{2}(G^{2}+\sigma_{l}^{2}).\end{split}$
Lemma 0 (Aggregated Gradient Second Moment Bound).
Under Assumptions 2, 3, and Lemma 2, assuming uniform client sampling on $n_{s}$ clients, we can derive the aggregated gradient second moment bound as followings:
(22) $\mathbb{E}[\|\bar{g}_{g}^{t}\|^{2}]\leq 2\mathbb{E}[\|\nabla F(\Theta_{g}^{t})\|^{2}]+2\mathbb{E}[\|B_{t}\|^{2}]+\frac{2(\sigma^{2}+G^{2})}{n_{s}}$
Proof of Lemma 1 (Aggregated Gradient Second Moment).
When Assumptions 2 and 3 hold, we can use the variance decomposition to derive the followings:
(23) $\begin{split}\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}&=\Bigl{\|}\mathbb{E}\bar{g}_{g}^{t}\Bigr{\|}^{2}+\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}-\mathbb{E}\bar{g}_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &=\Bigl{\|}\nabla F(\Theta_{g}^{t})+B_{t}\Bigr{\|}^{2}+\operatorname{Var}\left(\bar{g}_{g}^{t}\right)\\[3.0pt] &\leq 2\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+2\|B_{t}\|^{2}+\frac{1}{n_{s}}\,\mathbb{E}_{u}\Bigl{\|}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}\Bigr{\|}^{2}\\[3.0pt] &\leq 2\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+2\|B_{t}\|^{2}+\frac{G^{2}+\sigma^{2}}{n_{s}}.\end{split}$
The last step is based on Assumptions 2 and 3. ∎
Proof of Theorem 1 (Convergence Rate).
Starting from the $L$-smoothness of $F$, which is in Assumpution 1, we have:
(24) $\begin{split}F(\Theta_{g}^{t+1})&\leq F(\Theta_{g}^{t})+\bigl{\langle}\nabla F(\Theta_{g}^{t}),\Theta_{g}^{t+1}-\Theta_{g}^{t}\bigr{\rangle}+\frac{L}{2}\Bigl{\|}\Theta_{g}^{t+1}-\Theta_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &=F(\Theta_{g}^{t})-\eta\,\bigl{\langle}\nabla F(\Theta_{g}^{t}),\bar{g}_{g}^{t}\bigr{\rangle}+\frac{L\eta^{2}}{2}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}.\end{split}$
Taking expectations and using Lemmas 2 and 3, we obtain:
(25) $\begin{split}\mathbb{E}F(\Theta_{g}^{t+1})&\leq\mathbb{E}F(\Theta_{g}^{t})-\eta\,\mathbb{E}\Bigl{\langle}\nabla F(\Theta_{g}^{t}),\nabla F(\Theta_{g}^{t})+B_{t}\Bigr{\rangle}+\frac{L\eta^{2}}{2}\,\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &\leq\mathbb{E}F(\Theta_{g}^{t})-\frac{\eta}{2}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+\frac{\eta}{2}\,\mathbb{E}\|B_{t}\|^{2}\\[3.0pt] &\quad+L\eta^{2}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+L\eta^{2}\,\mathbb{E}\|B_{t}\|^{2}+\frac{L\eta^{2}(G^{2}+\sigma^{2})}{2n_{s}}.\end{split}$
Here we see the interplay of the gradient norm $\|\nabla F(\Theta_{g}^{t})\|$ and the bias term $B_{t}$. We control these via the learning rate $\eta$. Choosing $\eta\leq\frac{1}{4L}$ ensures certain terms are upper-bounded by $\frac{1}{2}\|\nabla F(\Theta_{g}^{t})\|^{2}$.
For $\eta\leq\frac{1}{4L}$, collecting like terms yields:
(26) $\begin{split}\frac{\eta}{4}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}&\leq\mathbb{E}\Bigl{[}F(\Theta_{g}^{t})-F(\Theta_{g}^{t+1})\Bigr{]}+\frac{5L^{3}C_{1}A\,\eta^{3}(G^{2}+\sigma_{l}^{2})}{2}\\ &\quad+\frac{L\eta^{2}(G^{2}+\sigma^{2})}{2n_{s}}.\end{split}$
Summing over $t=0$ to $T-1$ and setting $\eta=\frac{1}{L\sqrt{T}}$, we obtain
(27) $\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}&\leq\frac{4L\,(F_{0}-F^{*})}{\sqrt{T}}+\frac{10C_{1}AG^{2}}{T}+\frac{2(G^{2}+\sigma^{2})}{\sqrt{T}\,n_{s}}.\end{split}$
The additional term $\mathcal{O}\bigl{(}G^{2}/T\bigr{)}$ is dominated by $\mathcal{O}\bigl{(}G^{2}/\sqrt{T}\bigr{)}$, which concludes the proof. ∎
@@ B.1. Convergence Analysis
Appendix C Parameter Analysis
@@
Despite the increase in the number of learnable parameters, Table 2 of the main paper demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset. By combining multimodal features with ID embeddings, FedMR better leverages users’ interaction histories to capture the fine-grained personality of users. Additionally, the mixing feature fusion module in FedMR adjusts the weights of multimodal features based on users’ interaction history, greatly enhancing both the accuracy and personalization of recommendations. Moreover, compared to traditional FedRec methods, which rely solely on the ID embeddings, FedMR enables personalized modeling by integrating multimodal information, thereby further improving recommendation quality. While the model’s complexity increases, this trade-off results in significant performance gains.
@@
Appendix D More Experiments
D.1. More Experiment Details
All baseline methods are implemented in PyTorch (Paszke et al., 2019), and their official implementations are publicly available. Experiments are conducted on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, three Quadro RTX 6000 GPUs, and 395 GB of RAM.
Figure 4. Training loss progression over epochs for baseline methods enhanced with FedMR on the KU dataset, illustrating convergence behavior.
D.1.1. Convergence Analysis
To assess the training stability and convergence properties of the baseline methods when enhanced with FedMR, we plot their training loss progression over 50 epochs on the KU dataset in Figure 4. The loss curves demonstrate that all FedMR-enhanced models exhibit stable convergence behavior. The training loss decreases consistently across epochs for all methods, with a sharp reduction typically observed in the initial epochs followed by more gradual convergence towards a stable value. This indicates that the FedMR framework integrates well with diverse base FedRec architectures, allowing for effective optimization and stable learning dynamics. While the final converged loss values vary slightly across methods, all models successfully minimize the training objective, confirming the robustness of the training process when incorporating FedMR’s multimodal fusion capabilities.
Figure 5. Sensitivity analysis of 4 FedMR-enhanced baselines (FCF, FedAvg, FedNCF, PFedRec) on the KU dataset. Performance is plotted against varying learning rates (lr)and L2 regularization (l2_reg) strengths.
Figure 6. Hyperparameter tuning for FedMR-enhanced FedRAP on the KU dataset. Heatmaps visualize HR@50 and NDCG@50 across personalization parameters $\alpha$ and $\beta$ for different learning rates (lr) and L2 regularization (l2_reg). Performance intensity is indicated by color.
D.1.2. Hyperparameter Analysis
We analyze the sensitivity of FedMR-enhanced models to key hyperparameters, specifically the learning rate and L2 regularization strength, using the KU dataset. Fig. 5 illustrates this for FedMR applied to FCF, FedAvg, FedNCF, and PFedRec. Across these models, the learning rate exhibits a significant impact, with performance generally peaking at intermediate values and decreasing notably at lower or higher rates. This highlights the necessity of selecting an appropriate learning rate for effective convergence. The influence of L2 regularization appears moderate in comparison, though still relevant; no single $l2\_reg$ value is universally optimal, with the best performance achieved at different strengths depending on the base model and learning rate, necessitating model-specific tuning.
The tuning process for the FedMR-enhanced FedRAP model, which includes its specific personalization parameters $\alpha$ and $\beta$, is detailed in Fig. 6. The heatmaps clearly indicate high sensitivity to the choices of $\alpha$ and $\beta$, with distinct optimal regions emerging. Furthermore, these optimal regions shift based on the learning rate, signifying a strong interaction between $lr$, $\alpha$, and $\beta$. This underscores the need for careful co-tuning, typically via grid search, to identify the best configuration for FedMR-enhanced FedRAP. Overall, while the models demonstrate reasonable robustness, achieving optimal results requires careful hyperparameter tuning, particularly for the learning rate and any model-specific parameters. The best hyperparameters identified through validation during this analysis informed the settings used for the main experimental results.
D.1.3. Ablation Study on Modality Distributions
To analyze the interaction between multimodal features and ID features, we use FedRAP as the backbone for FedMR. Fig. 7 visualizes the distribution of these features on the KU dataset. The first subfigure shows that globally shared ID features and multimodal features occupy distinct subregions in the shared representation space, a phenomenon known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a). This separation enhances the robustness of multimodal data processing by reducing noise interference. The subsequent subfigures illustrate how personalized ID features influence this distribution. Incorporating personalized ID embeddings (e.g., for Users 00 and $42$) causes the ID features to expand toward the regions of visual and textual features, effectively shrinking the modality gap. This adjustment facilitates multimodal fusion and improves the accuracy of personalized recommendations. By comparing users, it is evident that the model dynamically adapts feature distributions based on user-specific behaviors and preferences, better addressing individual needs. This demonstrates the critical role of personalized ID features in federated recommendation systems, enabling effective personalization while preserving user privacy.
@@ D.1.3. Ablation Study on Modality Distributions
D.2. Ablation Study on Client Sampling
@@
Table 6. Performance comparison on the KU dataset. The table reports HR@50 and NDCG@50 for several FedRecs. Percentages indicate the performance drop relative to the full-feature baseline.
Method Metric FedMR w/o V w/ rnd V w/o C w/ rnd C w/o D w/ rnd D
FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3922 (23.80% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.3676 (28.58% $\downarrow$) 0.5094 (1.03% $\downarrow$) 0.4520 (12.18% $\downarrow$)
NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1303 (12.61% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1240 (16.83% $\downarrow$) 0.1407 (5.66% $\downarrow$) 0.1375 (7.78% $\downarrow$)
FCF HR@50 0.5539 0.2843 (48.67% $\downarrow$) 0.2892 (47.79% $\downarrow$) 0.4216 (23.89% $\downarrow$) 0.2451 (55.75% $\downarrow$) 0.2353 (57.52% $\downarrow$) 0.2206 (60.17% $\downarrow$)
NDCG@50 0.1597 0.0748 (53.16% $\downarrow$) 0.0783 (50.97% $\downarrow$) 0.0965 (39.57% $\downarrow$) 0.0629 (60.61% $\downarrow$) 0.0578 (63.81% $\downarrow$) 0.0538 (66.31% $\downarrow$)
FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1667 (14.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1618 (17.49% $\downarrow$) 0.1602 (18.31% $\downarrow$) 0.1569 (19.99% $\downarrow$)
NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0489 (19.70% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0461 (24.30% $\downarrow$) 0.0538 (11.66% $\downarrow$) 0.0458 (24.79% $\downarrow$)
PFedRec HR@50 0.2157 0.1990 (7.74% $\downarrow$) 0.1647 (23.64% $\downarrow$) 0.1941 (10.01% $\downarrow$) 0.1696 (21.37% $\downarrow$) 0.1794 (16.83% $\downarrow$) 0.1647 (23.64% $\downarrow$)
NDCG@50 0.0616 0.0469 (23.86% $\downarrow$) 0.0468 (24.03% $\downarrow$) 0.0537 (12.82% $\downarrow$) 0.0445 (27.76% $\downarrow$) 0.0473 (23.21% $\downarrow$) 0.0484 (21.43% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2206 (42.31% $\downarrow$) 0.3235 (15.40% $\downarrow$) 0.2304 (39.75% $\downarrow$) 0.1863 (51.29% $\downarrow$) 0.1324 (65.38% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0634 (51.42% $\downarrow$) 0.0714 (45.29% $\downarrow$) 0.0564 (56.78% $\downarrow$) 0.0490 (62.43% $\downarrow$) 0.0338 (74.10% $\downarrow$)
D.2.1. Ablation Study on Modality Contribution
We present the full modality contribution analysis, including results for noise replacement ($w/rnd$), in Table 6 owing to page limits.
We assess the individual contributions of visual ($\mathbf{V}$), textual ($\mathbf{C}$), and ID ($\mathbf{D}$) feature types within FedMR by evaluating performance upon either removing the feature type entirely ($w/o$) or replacing it with random noise ($w/rnd$). Table 6 presents these results on the KU dataset, revealing the distinct and complementary roles of each modality. The comparison between feature removal and noise replacement elucidates the nature of each modality’s contribution.
ID features consistently prove indispensable for personalization. Their removal ($w/o~{}D$) incurs substantial performance degradation across all methods. Notably, replacing ID features with random noise ($w/rnd~{}D$) often leads to even more severe performance drops, particularly for models with strong personalization components like FCF and FedRAP. This trend underscores that models heavily rely not just on the presence of an ID signal, but on the specific collaborative and structural information learned within the ID embeddings.
Visual features also demonstrate significant influence. Removing them ($w/o~{}V$) markedly hinders performance, especially for architectures known to leverage visual cues, such as FCF and FedRAP. Comparing this to noise replacement ($w/rnd~{}V$), we observe that substituting visual features with noise frequently results in equal or greater performance degradation than simple removal. This finding suggests that visual features provide meaningful, non-random signals that benefit recommendation, and replacing them with noise is often more detrimental than their absence, particularly for visually sensitive models.
Textual features provide valuable complementary semantic context. While their removal ($w/o~{}C$) impacts performance to a lesser extent than ID or visual features for some models, the effect is still notable across various methods. Replacing textual features with noise ($w/rnd~{}C$) generally causes a more significant performance decrease than simple removal, confirming that the semantic content captured by text offers a non-trivial contribution beyond just acting as a placeholder.
Collectively, these results highlight a clear interdependence among ID, visual, and textual information sources. The consistent performance drop upon removal or noise replacement confirms that each modality contributes uniquely. The often more detrimental effect of noise replacement compared to feature removal strongly indicates that the learned multimodal features contain meaningful signals crucial for performance. Furthermore, the degree of reliance on each modality, and the differential impact of removal versus noise replacement, varies markedly across the underlying FedRec architectures. This complex interplay and differential dependency motivate the need for adaptive fusion mechanisms. FedMR’s MFFM addresses this by dynamically balancing the contributions from these distinct and valuable information sources based on user context, enabling robust performance through effective information leverage.
Appendix E More Related Work
E.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024b), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024c) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
E.2. Multimodal Recommendation Systems
E.3. Foundation Models
Appendix F Limitations and Future Work
While FedMR demonstrates promising results for privacy-preserving multimodal recommendation, key limitations suggest avenues for future research. The specific subfield of FedMMRec is in its early stages: while FedMR advances this area alongside recent works like AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a), establishing direct experimental comparisons to precisely gauge relative performance remains a key challenge for future work as the field evolves and common benchmarks emerge. Additionally, FedMR primarily utilizes visual and textual data via CLIP, future research should explore incorporating a broader range of relevant modalities, such as audio or temporal information from videos. Further investigations could also focus on enhancing the client-side MFFM personalization mechanism or integrating FedMR with advanced FL techniques to better address communication constraints or statistical heterogeneity.